40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 23 23:42:56 2015, ericq, Update, LSC, ALSY recovered 
    Reply  Tue Nov 24 01:14:23 2015, Koji, Update, LSC, ALSY recovered 
Message ID: 11803     Entry time: Mon Nov 23 23:42:56 2015     Reply to this: 11804
Author: ericq 
Type: Update 
Category: LSC 
Subject: ALSY recovered 

[ericq, gautam]

Gautam couldn't observe a Y green beatnote earlier, so we checked things out, fixed things up, and performance is back to nominal based on past references. 

Things done:

  • Marconi carrier output switched back on after Koji's excellent RF maintence
  • BBPD power supplies switched on
  • Removed a steering mirror from the green beatY path to do near/far field alignment. 
  • Aligned PSL / Y green beams 
  • Replaced mirror, centered beam on BBPD, moved GTRY camera to get the new spot.
  • POY locked, dither aligned, beatnote found, checked ALS out-of-loop noise, found to be in good shape. 
ELOG V3.1.3-